# EraseBench: Understanding The Ripple Effects of Concept Erasure Techniques

### v1

@@ Abstract
Concept erasure techniques have recently gained significant attention for their potential to remove unwanted concepts from text-to-image models. While these methods often demonstrate success in controlled scenarios, their robustness in real-world applications and readiness for deployment remain uncertain. In this work, we identify a critical gap in evaluating sanitized models, particularly in terms of their performance across various concept dimensions. We systematically investigate the failure modes of current concept erasure techniques, with a focus on visually similar, binomial, and semantically related concepts. We propose that these interconnected relationships give rise to a phenomenon of concept entanglement resulting in ripple effects and degradation in image quality. To facilitate more comprehensive evaluation, we introduce EraseBENCH, a multi-dimensional benchmark designed to assess concept erasure methods with greater depth. Our dataset includes over 100 diverse concepts and more than 1,000 tailored prompts, paired with a comprehensive suite of metrics that together offer a holistic view of erasure efficacy. Our findings reveal that even state-of-the-art techniques struggle with maintaining quality post-erasure, indicating that these approaches are not yet ready for real-world deployment. This highlights the gap in reliability of the concept erasure techniques.
Figure 1: The ripple effects of concept erasure on ’entangled’ concepts. The images on the left show the output generated by Stable Diffusion (SD) prior to erasure, while those on the right depict the post-erasure results. We observe a noticeable negative impact on the generation quality of concept that are not erased. Assessment using our EraseBench framework helps identify these effects and provides a framework to evaluate the reliability and robustness of concept erasure techniques.
$\dagger$\$$\dagger$\$footnotetext:  The work was done during an internship at Google.Corresponding authors: ibtihel.amara@mail.mcgill.ca,
mhavaei@google.com, and nrostamzadeh@google.com
@@ 1 Introduction
As text-to-image generative models [33, 31, 16, 37, 6, 15] have become increasingly popular and widely adopted, the demand for techniques that allow for the targeted removal of specific concepts has grown in parallel. Concept erasure methods [27, 25, 7, 8, 13, 23, 47] promise to address important concerns such as ensuring the safe deployment of models by removing undesirable or harmful concepts [7, 10, 44]. Whether applied to mitigate bias [3, 38, 27], protect privacy [44, 17, 7, 10], or filter out unwanted content. These techniques aim to create ”sanitized” models capable of generating content without triggering unwanted associations.
Generative models interact with a variety of contexts and need to handle diverse, interrelated concepts. The ability to erase one concept without inadvertently affecting related or visually similar concepts remains a critical challenge. For example, erasing ”cat” from a model should ideally remove all representations of a cat, yet leave semantically or visually related concepts like ”lion” unaffected. Figure 1 shows various failures modes observed in state-of-the-art concept erasure techniques. These techniques are highly sensitive after a concept has been erased. Closely related non-target concepts often suffer unintended consequences, resulting in poor-quality image generation or even complete mode collapse. In this work, we argue that current concept erasure techniques often fall short in such real-world conditions. Specifically, we show that even state-of-the-art methods are vulnerable to failure modes when tasked with erasing concepts that are visually similar, binomially related, or semantically entangled. This lack of robustness indicates that these models may not be ready for deployment in sensitive applications and may have ripple effects that extend beyond their intended targets, potentially leading to unintended consequences such as the distortion of related concepts, which compromises the overall reliability of the text-to-image models.
To address this gap, we introduce EraseBENCH, a comprehensive benchmark designed to stress-test concept erasure techniques across multiple dimensions. By evaluating models on a wide range of concepts—spanning different domains, relationships, and complexities—EraseBENCH enables more realistic assessments of concept erasure performance. We present a novel dataset of over 100 diverse concepts and more than 1,000 tailored prompts, along with a suite of metrics to evaluate erasure efficacy, quality retention, and potential leakage of erased concepts. Through our evaluations, we reveal that many erasure techniques struggle with maintaining the generation quality and integrity post-erasure, underscoring the need for continued research in this space.
@@
    We present EraseBench, a comprehensive, multi-dimensional benchmark designed to rigorously evaluate the robustness and efficacy of concept erasure techniques across a diverse set of concepts and prompts.
@@
    We used a suite of evaluation metrics that captures the effectiveness of concept erasure methods, offering a more holistic assessment of their performance.
@@
    We benchmark five state-of-the-art concept erasure techniques on EraseBench and through this evaluation, demonstrate that many of the current concept erasure techniques are not yet ready for real-world deployment, highlighting significant gaps in their reliability and robustness.
@@ 2 Related Works
Text-to-image models have witnessed remarkable advancements since the foundational works [50, 24], driven by breakthroughs in model architectures [45, 49, 36, 41, 14, 5], state-of-the-art generative modeling techniques [30, 11, 31, 1, 33, 26, 2, 22, 19, 22, 9, 32, 18, 42], and the availability of large high-quality datasets [34, 20]. These improvements have significantly enhanced their capability to generate realistic and diverse images from textual descriptions, pushing the boundaries of creativity and application. In light of recent advancements in text-to-image models, greater attention has been directed towards reliability considerations. There is a wide variety of concept erasure techniques, each utilizing distinct methods and technologies to achieve concept removal. These approaches range from closed form solutions [8], applying targeted fine-tuning [7, 47, 13, 23], and even using efficient architectures such as adapters [13, 23]. We provide in the Appendix more details about existing concept erasure techniques. Nevertheless, despite the rapid increase of these concept erasure methods, only a few works have critically examined the inherent limitations and vulnerabilities of concept erasure approaches. In particular, the work by (Pham et al.) [28] explores a fundamental vulnerability, which is the potential for erased concepts to be recovered through specific prompting. Meanwhile, other works such as of (Zhang et al.) [48] focus on the efficacy of erasure through a variety of objects and styles to evaluate if the erased concepts have been effectively removed from the model’s generative capabilities. In our work, we delve into the broader impact of concept erasure, not only evaluating residual concept generation but also examining how erasure techniques may degrade the quality and fidelity of non-target concepts. In this work we highlight the ripple effects of concept erasure on model performance and the stability of related concepts via the route of interrelated concepts.
3 Concept Entanglement
Figure 2: Concept entanglement in SD. We use DAAM [35] to obtain attribution heatmaps for Stable Diffusion (SD) generations. While ”cat” and ”elephant” token attributions don’t overlap, for ”snake” and ”turtle” we can see both the snake and turtle objects being attributed, indicating an entanglement between concepts snake and turtle in SD. In Table. 1 we present four dimensions of entanglement between concepts and present empirical evidence of ripple effects post-erasure.
Our initial hypothesis stems from observing that many existing evaluations focused on isolated concepts, failing to account for the intricate relationships between concepts. Concept erasure techniques are designed to selectively remove specific concepts from generative models, such as text-to-image models, while maintaining the integrity of the remaining knowledge. However, this task becomes increasingly complex when dealing with concepts that are intertwined, either visually or semantically. Intertwined concepts refer to those that share overlapping features, either through visual similarity, semantic connection, or frequent co-occurrence in real-world data (i.e. binomial pairs).
The study of concept entanglement has long been a focus in the literature mainly focusing on exploring how to separate and understand intertwined concepts in representations for better image generation  [43, 40, 4, 46]. However, to our knowledge, these discussions have not yet been extended to the context of concept erasure.
Figure 2 showcases the attention attribution heatmaps [35]. For two closely related or co-occuring concepts in the prompt we observe significant overlap in the attention maps, with both concepts being activated by similar visual elements. In contrast, for two unrelated concepts—such as ’cat’ and ’elephant’ the attention maps remain well-separated. We argue that this entanglement results in unintended interference during erasure, making removing or isolating a concept significantly challenging.
In this work, we focus on evaluating the effects of entanglement within the context of concept erasure. Specifically, we analyze this entanglement under four lenses: (1) visual similarity, where concepts share visible characteristics (e.g., tigers and lions); (2) stylistic similarity in art, where visual elements convey similar aesthetics or themes; (3) binomiality, referring to culturally or popularly associated concept pairs (e.g., ’lock and key’); and (4) hierarchical relationships (i.e., subset of a superset), where one concept represents a specific instance within a broader category.
4 EraseBench
To evaluate the robustness of concept erasure techniques, we introduce a dataset designed to challenge model performance across several critical dimensions. Each dimension represents a unique relationship or similarity type that can impact the effectiveness and accuracy of concept erasure. By addressing these dimensions, the dataset provides a comprehensive foundation for understanding where current methods may falter, specifically in managing related and intertwined concepts.
4.1 Dimensions of evaluation
Visual Similarity in Objects: This category includes objects with overlapping visual features, such as cats, tigers, and lions. These classes are selected for their shared characteristics, which can lead models to generalize incorrectly, especially after a concept erasure operation. Testing against these visually similar objects helps reveal if erasure inadvertently impacts non-target but visually close concepts, indicating a lack of precision in the erasure.
Visual Similarity in Art: This dimension focuses on diverse artist signatures that share aesthetic or stylistic attributes or works by closely related artists. Since many erasure methods operate at the level of visual patterns, the subtle differences in style may not be adequately preserved, causing issues with fidelity in generation. This serves as a test for style and nuance preservation post-erasure.
Binomial Concepts: Binomial or paired concepts, like ”bread and butter,” present a unique challenge for erasure methods due to their strong associative links. This dimension examines how well models handle the erasure of one concept when it is closely linked with another. The goal is to identify if erasure leads to unintended consequences on the associated concept, revealing potential overreach in the erasure process.
Subset-Superset Relationships: This dimension explores hierarchical relationships or what we can call partial erasure, such as “black cat” within the broader “cat” category. Concept erasure techniques should ideally remove specific attributes or concepts while preserving general or encompassing ideas. Testing subset-superset relationships reveals whether erasure disrupts generalized concepts when only specific instances are targeted, offering insights into the granularity and control of the erasure process.
Erase: Mouse
Prompt: A hamster
Original Image
@@ 4.1 Dimensions of evaluation
Erase: Koala
Pr.: A tree kangaroo
Visual Similarity (Object)
@@
Visual Similarity (Art)
Erase: Latte
Prompt: Espresso
Erase: Lemon
Prompt: Orange
@@
4.2 Curating EraseBench
@@
Concept Gathering. We utilized the semantic knowledge embedded in large language models and structured sources, like the hierarchical taxonomy of ImageNet, to generate clusters of closely related concepts. By treating these language models akin to graph neural networks, we mapped visually similar concepts that share key attributes and gathered distinct sets for each dimension in EraseBench. This approach enabled the capture of nuanced similarities and relationships, essential for challenging the precision of concept erasure techniques. As for artistic style concepts, we narrowed down artists’ names from the Wikiart dataset.
@@
Concept Verification. To ensure quality and feasibility, we conducted human verification on the collected concepts. This included assessing whether existing text-to-image models could successfully generate accurate, high-quality representations of each concept. During this process, we identified and removed any concepts that models consistently struggled to render, preserving only those concepts that aligned with our benchmarks. This step ensures that the dataset maintains a high standard, with each concept suitable for rigorous testing of erasure techniques.
@@
Prompt Construction. We employed large language models to generate a diverse set of prompts for each concept. Prompts were crafted with variations in style, complexity, and length to allow for thorough evaluation. This includes prompts that differ in level of detail (simple vs. elaborate descriptions) and length (short vs. extended prompts). With this range, we enable a finer assessment of how concept erasure techniques handle varying input types and examine the degree of degradation under different conditions.
4.2.1 Concept Description
For each evaluation dimension, we define a primary concept intended for erasure, accompanied by related non-target concepts that share visual, semantic, or contextual similarities. These non-target concepts are used to prompt the sanitized model, allowing us to observe any unintended effects.
(1) Visual Similarity (Objects): We select a broad main concept, include four paraphrased (synonym) concepts, and add four visually similar concepts. For example, the main concept “cat” is erased, with synonyms like “kitten,” “tabby,” “siamese,” and “British shorthair,” while visually similar non-targets include “tiger,” “cheetah,” “lynx,” and “panther.” (2) Visual Similarity (Artist Style): Here, we erase a primary artist, e.g., Vincent van Gogh, and test model behavior on artists with similar styles, such as Paul Cézanne and Émile Bernard. (3) Subset-Superset Relationship: A primary target, such as “goldfish,” is erased, while testing is done on related types like “guppy” and “koi” to assess model retention of similar concepts within the category. (4) Binomial Relations: We use pairs of closely associated concepts, such as “sun and moon,” erasing one while testing generation for the other. Detailed descriptions of each concept set are available in the Appendix.
4.3 Baseline Methods
@@
Table 1: Evaluation of Erasing concepts from EraseBench under the four different dimensions. We provide the average results across 10 concepts (for visual similarity (object), across 15 artists concepts (for artists similarity), 8 concepts for subset-superset, and 10 concepts (for binomial concepts). CLIP’s zero-shot prediction along with the standard error mean (accuracies) are reported for each dimension in four sets: the erased concept (Eff., Efficacy), the paraphrase concepts (Gen., generality), the non-target visually similar concepts (Sens., sensitivity). The values presented are percentages (%). We provide more detailed results per concept in the Appendix. The classification accuracies of images generated by the original SD v1.4 are presented for reference.
@@
Visual Similarity (Object) Visual Similarity (Art) Subset & Super-set Binomial
Techniques Eff. $\downarrow$ Gen. $\downarrow$ Sens.$\uparrow$ Eff.$\downarrow$ Sens.$\uparrow$ Eff. $\downarrow$ Sens.$\uparrow$ Sens.$\uparrow$
Original 86.5 $\pm$ 7.1 90.2 $\pm$ 3.6 85.0 $\pm$ 4.0 72.3 $\pm$ 6.3 69.05 $\pm$ 4.3 83.3 $\pm$ 8.9 92.1 $\pm$ 3.5 88.5 ±$\pm 4.5$
ESD [7] 24.5 $\pm$ 6.1 50.52 $\pm$ 4.1 65.9 $\pm$ 4.6 15.1 $\pm$ 3.7 40.3 $\pm$ 3.9 18.1 $\pm$ 4.2 65.1 $\pm$ 5.1 70.6 $\pm$ 4.0
UCE [8] 41.8 $\pm$ 5.5 68.3 $\pm$ 2.7 82.7 $\pm$ 3.1 21.1 $\pm$ 4.3 61 $\pm$ 4.3 51.1 $\pm$ 8.8 87.5 $\pm$ 2.9 86.1 $\pm$4.32
Receler [13] 8.1 $\pm$ 3.2 20 $\pm$ 3.4 65.4 $\pm$ 3.8 8.7 $\pm$2.2 22.8 $\pm$3.8 4.2 $\pm$ 0.8 36.7 $\pm$ 7.5 57.5 $\pm$ 9.2
MACE [23] 15.6 $\pm$ 6.4 37.7 $\pm$ 3.9 66.4 $\pm$ 4.3 20.2 $\pm$ 4.6 49.2$\pm$ 4.8 13.9 $\pm$3.2 66.9 $\pm$ 5.6 70.7 $\pm$ 5.3
AdvUnlearn [47] 8.7 $\pm$ 2.9 39.1 $\pm$ 6.5 64.3 $\pm$ 5.7 14.5 $\pm$ 4.05 27.4 $\pm$ 4.0 7.4 $\pm$ 2.6 60.1 $\pm$ 5.9 64.6 $\pm$ 5.9
4.4 Experimental Setup

### v2

# Erasing More Than Intended? How Concept Erasure Degrades the Generation of Non-Target Concepts

@@ Abstract
Concept erasure techniques have recently gained significant attention for their potential to remove unwanted concepts from text-to-image models. While these methods often demonstrate promising results in controlled settings, their robustness in real-world applications and suitability for deployment remain uncertain. In this work, we (1) identify a critical gap in evaluating sanitized models, particularly in assessing their performance across diverse concept dimensions, and (2) systematically analyze the failure modes of text-to-image models post-erasure. We focus on the unintended consequences of concept removal on non-target concepts across different levels of interconnected relationships including visually similar, binomial, and semantically related concepts. To address this, we introduce EraseBench, a comprehensive benchmark for evaluating post-erasure performance. EraseBench includes over 100 curated concepts, targeted evaluation prompts, and a robust set of metrics to assess both effectiveness and side effects of erasure. Our findings reveal a phenomenon of concept entanglement, where erasure leads to unintended suppression of non-target concepts, causing spillover degradation that manifests as distortions and a decline in generation quality.
Figure 1: The effects of concept erasure on non-target concepts. Pre-erasure outputs (left) vs. post-erasure results (right) from Stable Diffusion (SD). Erasure negatively impacts the quality of unrelated concepts. EraseBench identifies such effects and offers a framework to evaluate the reliability of erasure methods.
@@ 1 Introduction
As text-to-image (T2I) generative models [40, 38, 21, 45, 6, 20] have become increasingly popular and widely adopted, the demand for techniques that allow for the targeted removal of specific concepts has grown in parallel. Concept erasure methods [34, 32, 10, 11, 18, 30, 54, 13, 26, 25, 44] promise to address important concerns such as ensuring the safe deployment of models by removing undesirable or harmful concepts [10, 15, 52]. Whether applied to mitigate bias [4, 46, 34], protect privacy [52, 22, 10, 15], or filter out unwanted content, these techniques aim to create “sanitized” models capable of generating content without triggering unwanted associations. However, the challenge of effectively erasing a concept without unintentionally distorting related or visually similar concepts remains unresolved. For example, erasing “cat” from a model should ideally remove all representations of a cat, yet leave semantically or visually related concepts like “lion” unaffected. Figure 1 shows various failure modes observed in state-of-the-art concept erasure techniques. These techniques exhibit instability after a concept is erased, often causing unintended consequences for closely related non-target concepts such as over-erasure and image artifacts (i.e. distortions). In this work, we argue that concept erasure techniques often struggle under real-world conditions. Even state-of-the-art methods fail when erasing concepts that are visually similar, binomially related, or semantically entangled. This lack of robustness raises concerns about their deployment in sensitive applications, as unintended distortions of related non-target concepts affects the reliability of T2I models.
While recent works assess erasure techniques using arbitrary concepts, a standardized framework for stress-testing across diverse concept relationships is still lacking. To address this, we introduce EraseBench, a unified benchmark with curated concepts and evaluations designed to rigorously test T2I models post-erasure. Covering binomial pairs, visually similar objects, and semantic relations. EraseBench enables consistent evaluation of erasure effectiveness, unintended distortions, and concept leakage. Our findings highlight persistent challenges and the need for more robust and standardized testing. We envision EraseBench as a foundation for improving comparability and advancing safer and more reliable erasure methods. Additionally, our results demonstrate that these sanitized models are not yet ready for deployment, highlighting the need for more robust and reliable erasure techniques that can operate effectively in complex, real-world scenarios.
@@
    We present EraseBench, a comprehensive, multi-dimensional framework designed to rigorously evaluate the robustness and efficacy of concept erasure techniques across a diverse set of concepts and prompts.
@@
    We use a suite of evaluation metrics to measure the effectiveness of concept erasure methods that provides a more holistic assessment of their performance.
@@
    We evaluate five state-of-the-art concept erasure techniques on EraseBench and demonstrate that many of the current concept erasure techniques are not yet ready for real-world deployment, highlighting significant gaps in their reliability and robustness.
@@ 2 Related Works
Concept Erasure. T2I models have witnessed remarkable advancements since the foundational works of  [57, 31], driven by breakthroughs in model architectures [53, 56, 42, 49, 19, 5], state-of-the-art generative modeling techniques [37, 16, 38, 1, 40, 33, 3, 29, 24, 29, 12, 39, 23, 50], and the availability of large high-quality datasets [41, 27]. These improvements have significantly enhanced their capability to generate realistic and diverse images from textual descriptions, pushing the boundaries of creativity and application. In light of recent advancements in T2I models, greater attention has been directed towards reliability considerations. There is a wide variety of concept erasure techniques, each utilizing distinct methods and technologies to achieve concept removal [10, 11, 54, 18, 30, 30, 9, 14, 22, 7].
Challenges in Concept Erasure. Despite the rapid increase of these concept erasure methods, only a few works have critically examined the inherent limitations and vulnerabilities of concept erasure approaches. In particular, the work in  [35] explores a fundamental vulnerability, which is the potential for erased concepts to be recovered through specific prompting. Meanwhile, other works, such as in  [55], assess the effectiveness of concept erasure by examining a diverse range of objects and styles. Their approach evaluates whether the erased concepts have been completely removed from the model’s generative capabilities, ensuring that no traces remain in the generated outputs. In our work, we delve into the broader impact of concept erasure, not only evaluating residual concept generation but also examining how erasure techniques may degrade the quality and fidelity of non-target concepts. In this work, we highlight the ripple effects of concept erasure on model performance and the stability of related concepts via the route of interrelated concepts.
Erase: Koala
Pr.: A tree kangaroo
Original
@@
Visual Similarity
@@
Artistic Similarity
Erase:Ice-cream
Prompt: A Popsicle
@@
3 EraseBench
In this section, we provide a detailed breakdown of the process used to construct our benchmark.
@@ 3 EraseBench
Concept Gathering. We utilized the semantic knowledge embedded in large language models (LLMs) and structured sources, like the hierarchical taxonomy of ImageNet, to generate clusters of closely related concepts. By treating these language models akin to graph neural networks, we mapped similar concepts that share key attributes and gathered distinct sets for each dimension. This approach captures nuanced relationships crucial for testing erasure precision. For artistic styles, we selected artist names from the WikiArt dataset [47].
@@
Concept Verification. To ensure quality and feasibility, we conducted human verification on the collected concepts. This included assessing whether existing T2I models could successfully generate accurate, high-quality representations of each concept. During this process, we identified and removed any concepts that models consistently struggled to render, preserving only those concepts that aligned with our benchmarks. This step ensures that the dataset maintains a high standard, with each concept suitable for rigorous testing of erasure techniques.
@@
Prompt Construction. We used Gemini Flash [43], to generate a diverse set of prompts for each concept. Prompts were crafted with variations in style, complexity, and length to allow for thorough evaluation. This includes prompts that differ in level of detail (simple vs. elaborate descriptions) and length (short vs. extended prompts). With this range, we enable a finer assessment of concept erasure techniques.
Figure 3: Evaluation Dimensions of EraseBench.
Concept Dimensions. We introduce multiple evaluation dimensions, each capturing a distinct type of concept relationship or similarity, providing a structured way to evaluate the impact of erasure. Within each dimension, we provide a diverse set of concepts that challenge the model in different ways, ensuring a thorough evaluation. Figure 3 summarizes these dimensions in EraseBench. For each, we define a primary concept for erasure and related non-target concepts sharing visual, semantic, or contextual similarities to test unintended effects on the sanitized model. (1) Visual Similarity (Objects), A broad concept (e.g., “cat”) is erased, along with synonyms (“kitten,” “tabby”), while visually similar non-targets (“tiger,” “cheetah”) test unintended erasure. (2) Style Similarity (Artists), Erasing an artist (e.g., Van Gogh) tests model behavior on similar-style artists (e.g., Cézanne, Bernard). (3) Subset-Superset Relationship, Removing a concept (e.g., “goldfish”) evaluates retention of related non-targets (“guppy,” “koi”). (4) Binomial Relations, Erasing one of a closely linked pair (e.g., “sun” from “sun and moon”) tests impact on the other.
Explicit Content. Under each dimension of EraseBench, we also integrate explicit concepts to evaluate the handling of sensitive and safety-critical content (details can be found in Table 8 in the supplemental material). These include categories such as NSFW, culturally related, and broader safety-related themes. Incorporating these explicit concepts allows for a more rigorous assessment of how well erasure methods balance effectiveness, generation quality, and sensitivity.
@@
Table 1: Evaluation of Erasing concepts from EraseBench under the four different dimensions. We provide the average results across 10 concepts (for visual similarity (object)), 15 artists concepts (for artists similarity), 8 concepts for subset-superset, and 10 concepts (for binomial). We evaluate efficacy (Eff., capturing the efficacy of erasing the main concept), generality (Gen., capturing the impact on paraphrased or related concepts), and sensitivity (Sens., capturing the unintended effect on non-target yet related or similar concepts). All values are expressed as percentages (%). More detailed, per-concept breakdowns are provided in the Supplemental Material. For reference, we also include the classification accuracy of images generated by the original Stable Diffusion v1.4 model.
@@
Visual Similarity (Object) Artistic Style Similarity (Artists) Subset-Superset Binomial
[…]
Original 86.5 $\pm$ 7.1 90.2 $\pm$ 3.6 85.0 $\pm$ 4.0 15.97 72.3 $\pm$ 6.3 80.7 $\pm$ 3.2 69.0 $\pm$ 4.3 29.29 83.3 $\pm$ 8.9 92.8 $\pm$ 3.9 92.1 $\pm$ 3.5 14.31 88.3 $\pm$ 5.2 89.8 $\pm$ 4.3 88.5 $\pm 4.5$ 15.40
ESD [10] 24.5 $\pm$ 6.1 50.5 $\pm$ 4.1 65.9 $\pm$ 4.6 61.70 15.1 $\pm$ 3.7 61.9 $\pm$ 4.8 40.3 $\pm$ 3.9 47.74 18.1 $\pm$ 4.2 51.2 $\pm$ 4.2 65.1 $\pm$ 5.1 62.42 20.8 $\pm$ 4.6 41.3 $\pm$ 4.3 70.6 $\pm$ 4.0 68.46
UCE [11] 41.8 $\pm$ 5.5 68.3 $\pm$ 2.7 82.7 $\pm$ 3.1 49.32 21.1 $\pm$ 4.3 52.2 $\pm$ 3.5 61.0 $\pm$ 4.3 60.01 51.1 $\pm$ 8.8 62.7 $\pm$ 2.9 87.5 $\pm$ 2.9 51.12 18.9 $\pm$ 2.2 31.4 $\pm$ 2.8 86.1 $\pm$4.32 77.88
Receler [18] 8.1 $\pm$ 3.2 20 $\pm$ 3.4 65.4 $\pm$ 3.8 77.58 8.7 $\pm$2.2 45.3 $\pm$ 3.8 22.8 $\pm$3.8 41.04 4.2 $\pm$ 0.8 20.4 $\pm$ 3.7 36.7 $\pm$ 7.5 59.70 10.3 $\pm$ 2.4 12.6 $\pm$ 3.1 57.5 $\pm$ 9.2 75.04
[…]
AdvUnlearn [54] 8.7 $\pm$ 2.9 39.1 $\pm$ 6.5 64.3 $\pm$ 5.7 69.88 14.5 $\pm$ 4.1 37.4 $\pm$ 3.9 27.4 $\pm$ 4.0 47.75 7.4 $\pm$ 2.6 30.1 $\pm$ 5.2 60.1 $\pm$ 5.9 71.87 9.3 $\pm$ 2.2 27.6 $\pm$ 2.1 64.6 $\pm$ 5.9 74.41
4 Experimental setup
Baseline models. In our work, we narrowed down existing techniques and focused on leveraging more recent, advanced and diverse approaches to concept erasure, experimenting with recent techniques that incorporate variations such as fine-tuning model weights (e.g. ESD [10]), introducing targeted weight perturbations (e.g. UCE [11]), refining textual embeddings and adversarial training (e.g. Receler [18], AdvUnlearn [54]), and introducing parameter efficient fine-tuning (Receler, MACE [30]). This allows us to explore the nuanced dynamics of concept erasure and assess its impact under different dimensions. We provide detailed description about each of these techniques in the supplemental material.